OpenStudy (anonymous):

5x^2=10x A.(5,10) B.(0,5) C.(0,2) D.(2,5)

myininaya (myininaya):

are those suppose to be ordered pairs?

myininaya (myininaya):

do you want us to solve for x?

OpenStudy (anonymous):

5x^2=10x 5x^2-10x =0 5x(x-2)=0 x=0 or x=2 c.(0,2) is the correct choice